Need help on how to solve this problem. Please show steps on how to write the linear function...Thank You :)
Write the linear function for the real-world situation described.
Suppose that over a 20 year period, the number of residents of voting age increased at a nearly constant rate of 2.3 million people per year. After 11 years, there were 184.5 million residents of voting age. Model the number of residents of voting age (in millions) V as a function of the number of years y since the beginning of that 20 year period.

Answer by rothauserc(4718)

2.3 is our slope m of the line
we need to calculate the V intercept (V at year 0) which is
184.5 - (11*2.3) =
184.5 - 25.3 = 159.2
this means at year 0, V was 159.2 and therefore our equation is
V = 2.3y + 159.2
now we want to know V after 20 years
V = 2.3(20) + 159.2
V = 46 + 159.2
V = 205.2
After 20 years there are 205.2 million residents of voting age